Kay has amassed a collection of 1422 coins. they have included only nickels and pennies. the total collection is worth 45.22.
a. Write a system of equations to represent this information, and determine the number of each coin in the collection.

Answer by mananth(16946)

Kay has amassed a collection of 1422 coins. they have included only nickels and pennies. the total collection is worth 45.22.
a. Write a system of equations to represent this information, and determine the number of each coin in the collection.
a collection of 1422 coins.
nickels and pennies.
let the number of nickeis be x
and pennies be y


COIN count
x+y=1422..............................(1)
the total collection is worth 45.22.
Value count
5x+y=4522............................(2) (convert to pennies)
subtract eq 1 from 2
4x = 3100
x= 775 number of nickels
remaining pennies out of 1422